### v1

# Some Density Results for HyperKähler Manifolds

Abstract.
Lagrangian fibrations of hyperkähler manifolds are induced by semi-ample line bundles which are isotropic with respect to the Beauville-Bogomolov-Fujiki form. For a non-isotrivial family of hyperkähler manifolds over a complex manifold $S$ of positive dimension, we prove that the set of points in $S$, for which there is an isotropic class in the Picard lattice of the corresponding hyperkähler manifold represented as a fiber over that point, is analytically dense in $S$. We also prove the expected openness and density of the locus of polarised hyperkähler manifolds that admit a nef algebraic isotropic line bundle.
Key words and phrases:
Hyperkähler manifolds, Lagrangian fibration, Deformation, Density.
2020 Mathematics Subject Classification:
primary: 14D05, 14D06, 14J42. secondary: 37J38
Introduction
A hyperkähler manifold is a simply connected compact Kähler manifold $X$ of dimension $2n$ admitting a unique (up to scaling) global homolomorphic symplectic 2-form $\sigma\in H^{0}(X,\Omega_{X}^{2})$. The existence of such a 2-form makes $X$ Calabi–Yau in the sense that its canonical bundle is trivial. In dimension 2, these are K3 surfaces. In higher dimensions, two infinite series and two sporadic examples are known [Bea83, O’G99, O’G03], up to deformation.
Since any hyperkähler manifold $X$ comes equipped with a holomorphic symplectic form $\sigma$, Lagrangian subvarieties of $X$, i.e., subvarieties $Z\subset X$ of dimension $n=\frac{1}{2}\dim X$ such that $\sigma|_{Z^{\rm reg}}=0$, are interesting to study. It is known [Mat99] that any morphism $f\colon X\to B$ where $0<\dim B<2n$, is in fact a Lagrangian fibration, meaning the general fiber of $f$ is a Lagrangian subvariety. The general fibers are in fact abelian varieties, as proved by Voisin (see [Cam06, Prop. 2.1]). Moreover, whenever $B$ is smooth, it is in fact isomorphic to $\mathbb{P}^{n}$ [Hwa08]. Hence, such a Lagrangian fibration is induced by a semi-ample line bundle $\mathcal{L}$ on $X$ with $c_{1}(\mathcal{L})^{n+1}=0$.
The second cohomology $H^{2}(X,{\mathbb{Z}})$ comes equipped with a unique non-degenerate quadratic form $q_{X}$, the Beaville–Bogomolov–Fujiki form, which controls a great deal of geometric information about the hyperkähler manifold $X$. It has signature $(3,b_{2}-3),$ and there exists a non-zero constant $c_{X}\in\mathbb{Q}$ satisfying $\alpha^{2n}=c_{X}\cdot q_{X}(\alpha)^{n}$. In particular, admitting a semi-ample line bundle $\mathcal{L}$ with $c_{1}(\mathcal{L})^{n+1}=0$ is equivalent to $q_{X}(\mathcal{L})=0.$ Therefore, the first step in studying Lagrangian fibrations is the study of isotropic classes.
We prove the following:
Theorem 1.
Let ${\mathcal{X}}\rightarrow S$ be a non-isotrivial family of hyperkähler manifolds over a complex manifold $S$ with $\dim S>0.$ Then the set
$\{t\in S\mid\exists\alpha\in H^{1,1}({\mathcal{X}}_{t},\mathbb{Z}),q_{X}(\alpha)=0\}$
is dense in $S$ in the analytic topology.
The case when $S=\operatorname{Def}(X)$, the whole deformation space of $X$, is already known. For hyperkähler manifolds, there is a global Torelli theorem which allows one to prove such density statements directly on the period domain (i.e., the type IV symmetric domain associated to $(H^{2}(X,\mathbb{Z}),q_{X})$; see §1 for details). On the period domain, these types of density results follow from ergodic properties [Huy16, Prop. 7.1.3]. However, the question of density in an arbitrary family demands different methods. Our argument is Hodge theoretic and hence applies more generally to any non-isotrivial family of K3-type Hodge structures (see Theorem 1.2 for a precise statement). We follow [Voi07, §5.3.4].
One should compare Theorem 1.2 to the classical density result for Noether–Lefschetz loci, which, in the set-up of the theorem, states that the locus of hyperkähler manifolds where the Picard rank jumps is analytically dense in the base. The result is attributed to Oguiso [Ogu03, Theorem 1.1] and the key ingredient (also a key ingredient in our proof) is the criterion for the density of the Hodge classes, due to Green [Voi07, Proposition 5.20].
In the pursuit of Lagrangian fibrations, algebraic isotropic111isotropic = square zero classes as above are in a way insufficient; a hyperkähler manifold $X$ may have such an algebraic isotropic class $\alpha$, while none of the elements in its orbit under birational self-maps of $X$ are nef (see [Sac23, Thm. 4.1] for an example). Nefness is a necessary condition for such an $\alpha$ to induce a Lagrangian fibration. The hyperkähler SYZ conjecture postulates that it is in fact sufficient. The conjecture was first formulated by Strominger, Yau and Zaslow in [SYZ96] in the context of mirror symmetry of Calabi–Yau pairs. It has been reformulated for hyperkähler manifolds by Huybrechts, Hasset–Tschinkel, Tyurin, Bogomolov, and Sawon.
Conjecture 2.
[Saw03, Conj. 4.1, 4.2]
  1. (1)
    A hyperkähler manifold $X$ has a Lagrangian fibration if and only if it admits a non-trivial isotropic nef line bundle $L$.
  2. (2)
    A hyperkähler manifold $X$ admits a rational Lagrangian fibration (i.e., it is birational to a hyperkähler manifold which admits a Lagrangian fibration) if and only if it admits an isotropic non-trivial line bundle $L$.
Note that, according to [KL22, Lemma 2.18], for a fixed hyperkähler manifold $X$, then (1) is satisfied for all hyperkählers deformation equivalent to $X$ if and only if (2) is satisfied for all hyperkählers deformation to $X$. Therefore, one may focus on isotropic nef classes. To this end, Matsushita and Kamenova–Verbitsky studied the locus $\operatorname{Def}(X,L)_{\operatorname{reg}}$ of hyperkähler manifolds which admit a regular Lagrangian fibration in the deformation space $\operatorname{Def}(X,L),$ where $L$ is an algebraic isotropic line bundle. They showed:
Theorem 3.
[KV14, Theorem 3.4],[Mat17, Theorem 1.2] Let $X$ be a hyperkähler manifold and $L$ an isotropic line bundle on $X$. If $L$ induces a rational Lagrangian fibration, then $\operatorname{Def}(X,L)_{\operatorname{reg}}$ is a dense open subset of $\operatorname{Def}(X,L)$ and the set $\operatorname{Def}(X,L)\setminus\operatorname{Def}(X,L)_{\operatorname{reg}}$ is contained in a countable union of hypersurfaces in $\operatorname{Def}(X,L)$.
In light of the SYZ conjecture, it is expected that the locus $\operatorname{Def}(X,L)_{\operatorname{nef}}$ of hyperkähler manifolds which admit an algebraic isotropic nef class is also open and dense in $\operatorname{Def}(X,L)$. The hyperkähler SYZ conjecture is now established for all known deformation classes of hyperkähler manifolds. This is due to [BM14, Theorem 1.5] and [Mar14, Theorem 1.3, 6.3] for deformations of Hilbert schemes of points on a $K3$ surface, to [Yos16, Prop. 3.38] for deformations of generalized Kummer manifolds, [MR21, Cor. 1.3, 7.3] for deformations of $OG6$, and [MO22, Thm 2.2] for deformations of $OG10$. In the case of general polarised hyperkähler manifolds, we show this expected openness:
Theorem 4.
Let $X$ be a hyperkähler manifold and $L$ an isotropic nef line bundle on $X$. The locus $\operatorname{Def}(X,L)\setminus\operatorname{Def}(X,L)_{\rm nef}$ is a countable union of hypersurfaces in $\operatorname{Def}(X,L).$ Moreover, if, in addition, $X$ is a polarised hyperkähler manifold with ample line bundle $H$, then $\operatorname{Def}(X,L,H)_{\rm nef}$ is open and dense in $\operatorname{Def}(X,L,H)$.
The countable union of hypersurfaces in the complement of the locus $\operatorname{Def}(X,L)_{\rm nef}$, as mentioned in Theorem 4, is in fact induced by a countable union of the so-called wall divisors (see Definition 2.2). We use the fact that on the polarised period domain, the union of all hypersurfaces induced by these wall divisors is closed. This was shown by Amerik and Verbitsky. Roughly speaking, the closedness follows from the properly discontinous action of the group of isometries of the lattice on the period domain of polarised marked hyperkähler manifolds. Our proof critically uses this - see §2.
Let us finally note that questions about openness of the nef locus have been addressed in the literature beyond the hyperkähler setting [Mor92, Theorem 1.1]. As a consequence of Theorem 4 above, given a family of suitably polarised hyperkähler manifolds with an isotropic line bundle $L$ on it, Moriwaki’s result [Mor92] gives a bound on the minimum degree of the wall divisor that makes $L$ non-nef (see Corollary 2.5 for a precise statement).
1. Density of isotropic classes
Throughout, we denote by $(S,{\mathbb{V}},{\mathcal{F}}^{\bullet},\nabla,q)$ a polarised variation of Hodge structures (PVHS) over a connected complex manifold $S$. Here ${\mathbb{V}}$ is a local system of finite dimensional ${\mathbb{Q}}$-vector spaces with a Hodge filtration ${\mathcal{F}}^{\bullet}$ of the holomorphic bundle ${\mathcal{V}}:={\mathbb{V}}\otimes{\mathcal{O}}_{S}$, and $\nabla\colon{\mathcal{V}}\rightarrow{\mathcal{V}}\otimes\Omega_{S}^{1}$ is the Gauss-Manin connection. The polarisation $q$ is a flat bilinear form ${\mathbb{V}}\otimes{\mathbb{V}}\rightarrow{\mathbb{V}}$ whose restriction to each fiber $V_{s}$ of ${\mathbb{V}}$ at a point $s\in S$ is a polarisation of the Hodge structure on $V_{s}$.
Definition 1.1.
A finite dimensional vector space $V_{\mathbb{Q}}$ with a quadratic form $q$ is of K3-type if the signature of $q$ is $(3,\dim V_{\mathbb{R}}-3)$ with $\dim V_{\mathbb{Q}}\geq 4$. We say that a PVHS $(S,{\mathbb{V}},{\mathcal{F}}^{\bullet},\nabla,q)$ is of $K3$-type if it is a variation of polarised weight 2 Hodge structures with $\operatorname{rank}\mathcal{F}^{2}=1$ and each fiber is a K3-type vector space.
Given a K3-type vector space $(V_{\mathbb{Q}},q)$, the period domain of K3-type Hodge structures on $(V_{\mathbb{Q}},q)$ is defined as
$\Omega\coloneqq\{\sigma\in\mathbb{P}(V_{\mathbb{C}})\mid q(\sigma,\overline{\sigma})=0,q(\sigma)>0\}.$
Let $(S,{\mathbb{V}},{\mathcal{F}}^{\bullet},\nabla,q)$ be a PVHS of K3-type with a trivialisation $\varphi\colon\mathbb{V}\overset{\sim}{\to}V_{\mathbb{Q}}\otimes\mathbb{Q}_{S}$. We have the holomorphic period map:
$\mathcal{P}\colon S\to\Omega$
defined by $t\mapsto\varphi_{t}(\mathcal{F}^{2}_{t})$. We say that a PVHS of K3-type is non-isotrivial if the holomorphic period map $\mathcal{P}$ is non-constant.
For a class $\alpha\in V_{\mathbb{Q}}$, we denote by $\alpha^{\perp}(S)\subset S$ the locus:
$\alpha^{\perp}(S):=\{t\in S\mid\alpha\in\mathcal{F}^{1}_{t}\}.$
Note that $\alpha^{\perp}(S)=\{t\in S\mid\alpha\in V^{1,1}_{t}\}$, since $\alpha$ is equal to its complex conjugate.
We let $Q:=\{x\mid q(x)=0\}\subset V_{\mathbb{C}}$ be the quadric cone defined by the polarisation; we will assume that $Q$ has a nonzero rational point. Note that, if $\operatorname{rank}({\mathbb{V}})\geq 5$, this is automatically satisfied by Meyer’s theorem.
Theorem 1.2.
Let $(S,{\mathbb{V}},{\mathcal{F}}^{\bullet},\nabla,q)$ be a non-isotrivial PVHS of $K3$ type with trivialisation $\varphi\colon{\mathbb{V}}\overset{\sim}{\to}V_{\mathbb{Q}}\otimes{\mathbb{Q}}_{S}$ as above. Assume that the quadric cone $Q$ has a nonzero rational point. Then the union
$S_{\rm isot}:=\bigcup_{\begin{subarray}{c}\alpha\in V_{\mathbb{Q}}\\ q(\alpha)=0\end{subarray}}\alpha^{\perp}(S)$
is dense in $S$ in the analytic topology.
Proof.
Consider the subset
$Q^{1}:=\{(\lambda,s)\mid\lambda\in\mathcal{F}^{1}_{s},\,q(\lambda)=0\}\subset Q\times S$
and denote $p\colon Q^{1}\rightarrow Q$ the restriction of the projection to the first factor. By Lemma 1.3 below, the map $p$ is open around some point $(\lambda,s)\in Q^{1}$. Let $\pi\colon Q^{1}\to S$ be the restriction of the second projection. For any $s\in S$, there is an open neighbourhood $U\ni s$ such that $\pi^{-1}(U)\subset Q^{1}$ contains an open neighbourhood $V$ of $(\lambda,s)$ surjecting onto $U$ where $p|_{V}\colon V\to Q$ is open.
Note that the induced map on the real points $p_{\mathbb{R}}\colon V(\mathbb{R})\to Q(\mathbb{R})$ is also open. The sets $Q({\mathbb{Q}})$ and $Q({\mathbb{R}})$ are non-empty by assumption. Therefore, since $Q({\mathbb{Q}})\subset Q({\mathbb{R}})$ is dense, it follows that the subset $\{(\alpha,s)\mid\alpha\in\mathcal{F}^{1}_{s},q(\alpha)=0\}\cap V\subset V$ is dense. Therefore its image $\bigcup_{q(\alpha)=0}\alpha^{\perp}(U)$ in $U$ is also dense. ∎
It remains to show that the map $p:Q^{1}\rightarrow Q$ is a submersion. We do so below: we adapt the proof of [Voi07, Lemma 5.21] to our specific situation.
Lemma 1.3.
For any $s\in S$, there exists $\lambda\in Q^{1}_{s}$ such that the map $p\colon Q^{1}\rightarrow Q$ is a submersion at $(\lambda,s)$.
Proof.
Denote by $\widetilde{p}\colon\mathcal{F}^{1}\longrightarrow V_{\mathbb{C}}$ the projection to the first factor. By [Voi07, Lemma 5.21],
the map $\widetilde{p}$ is a submersion at $(\lambda,s)\in\mathcal{F}^{1}$ if the map
¯$\overline{\nabla}(\lambda)\colon T_{S,s}\rightarrow V_{s}^{0,2}$
induced by the Gauss–Manin connection is surjective. The map $\overline{\nabla}$ is also the differential $d\mathcal{P}$ of the period map ${\mathcal{P}}\colon S\rightarrow{\mathbb{P}}(V_{\mathbb{C}})$. This is because
¯$\overline{\nabla}\colon T_{S,s}\longrightarrow\operatorname{Hom}(V_{s}^{1,1},V_{s}^{0,2})$
and we have isomorphisms
$\operatorname{Hom}(V_{s}^{1,1},V_{s}^{0,2})\simeq\operatorname{Hom}((V_{s}^{0,2})^{*},(V_{s}^{1,1})^{*})\simeq\operatorname{Hom}(V_{s}^{2,0},V_{s}^{1,1})\simeq T_{[V_{s}^{2,0}]}\Omega.$
Since the PVHS $(S,{\mathbb{V}},\mathcal{F}^{\bullet},\nabla,q)$ is not isotrivial, this differential is not zero and hence there exists $v\in T_{S,s}$ such that $d\mathcal{P}(v)\colon V_{s}^{1,1}\to V^{0,2}_{s}$ is non-trivial, hence surjective. Therefore there exists $\lambda\in V_{s}^{1,1}$ such that ¯$\overline{\nabla}(\lambda)$ is surjective.
For a fixed $s$, the quadric $Q^{1}_{s}\cap V^{1,1}_{s}\subset V^{1,1}_{s}$ is not contained in a hyperplane by the definition of a polarisation. Since the dimension of $V_{s}^{1,1}$ is at least 2, $Q^{1}_{s}\cap V^{1,1}_{s}$ cannot be contained in the linear subspace $\operatorname{Ker}(d\mathcal{P}(v))\subset V_{s}^{1,1}$. Therefore, the map $\widetilde{p}$ is a submersion at $\lambda\in Q^{1}_{s}$.
Since ${Q^{1}}=\widetilde{p}^{-1}(Q)$, we obtain that for any $s\in S$, there exists $\lambda$ such that $p$ is a submersion at $(\lambda,s)$. ∎
We can apply Theorem 1.2 to non-isotrivial families of hyperkähler manifolds to recover Theorem 1:
Corollary 1.4.
[Theorem 1] Let ${\mathcal{X}}\rightarrow S$ be a non-isotrivial family of hyperkähler manifolds over a complex manifold $S$ with $\dim S>0.$ Then the union
$S_{\rm isot}:=\bigcup_{q(\alpha)=0}\alpha^{\perp}(S)$
is dense in $S$ in the analytic topology.
Remark 1.5.
One could also apply Theorem 1.2 to the Fano variety $F(X)$ of lines associated to a cubic 4-fold $X$. The Fano variety $F(X)$ is a projective hyperkähler manifold of K3${}^{[2]}$-type. By [Huy23, Corollary 5.24], if $F(X)$ admits a rational Lagrangian fibration, then the cubic fourfold $X$ lies on a special divisor $\mathscr{C}_{d}$ of the moduli of cubic fourfolds, indexed by an integer $d$ where $d/2$ is a perfect square. Said differently, Theorem 1.2 implies that any non-isotrivial family of cubic fourfolds contains a dense subset of points from $\mathscr{C}_{d}$ where $d/2$ is a perfect square.
Question 5.
In the light of the SYZ conjecture (see Conjecture 2), one may wonder whether, given a non-isotrivial family of hyperkähler manifolds ${\mathcal{X}}\to S$, the locus of algebraic isotropic nef classes is also dense.
2. Density of nef classes and Lagrangian fibrations
Throughout this section, we consider a marked hyperkähler manifold $X$ of a fixed deformation type. We denote by $\varphi:(H^{2}(X,{\mathbb{Z}}),q_{X})\overset{\sim}{\to}\Lambda$ the marking, i.e., $\varphi$ is an isometry of integral lattices. Let $L\in H^{1,1}(X,{\mathbb{Z}})$ be an isotropic (i.e., $q_{X}(L)=0$) primitive line bundle and let $\operatorname{Def}(X,L)$ be the hypersurface where $L$ remains algebraic [^2] Here we really mean the locus of $t\in\operatorname{Def}(X)$ where the class $\varphi_{t}^{-1}(\varphi(L))$ remains algebraic. in the local Kuranishi space $\operatorname{Def}(X)$. Let $({\mathcal{X}},{\mathcal{L}})\rightarrow\operatorname{Def}(X,L)$ be the universal family, together with a line bundle ${\mathcal{L}}$ such that $({\mathcal{X}},{\mathcal{L}})_{0}=(X,L)$. Note that the marking $\varphi$ of $X$ naturally defines markings $\varphi_{t}$ of the fibers ${\mathcal{X}}_{t}$, and thus we can consider the period map
${\mathcal{P}}\colon\operatorname{Def}(X)\rightarrow\Omega,$
where $\Omega:=\{\sigma\in{\mathbb{P}}(\Lambda\otimes{\mathbb{C}})\mid(\sigma\cdot\overline{\sigma})=0,\sigma^{2}>0\},$ as in §1. It follows that:
$\operatorname{Def}(X,L)={\mathcal{P}}^{-1}(\varphi(L)^{\perp}),$
with the notation of §1.
We will also consider the polarised case: if, in addition, $X$ admits an ample class $H$, we consider the universal family $({\mathcal{X}},{\mathcal{L}},\mathcal{H})\rightarrow\operatorname{Def}(X,L,H)$, together with an ample line bundle $\mathcal{H}$ such that $({\mathcal{X}},{\mathcal{L}},\mathcal{H})_{0}=(X,L,H).$ Note that
$\operatorname{Def}(X,L,H):=\operatorname{Def}(X,L)\cap\operatorname{Def}(X,H)\subset\operatorname{Def}(X);$
this is the locus $t\in\operatorname{Def}(X)$ where the classes $L$ and $H$ remain algebraic.
The aim of this section is to study the locus $\operatorname{Def}(X,L)_{\rm nef}\subset\operatorname{Def}(X,L)$ of points $t\in\operatorname{Def}(X,L)$ where the isotrivial primitive line bundle ${\mathcal{L}}_{t}$ is nef. According to the SYZ Conjecture 2, this is exactly the locus where ${\mathcal{L}}_{t}$ defines a Lagrangian fibration on $X_{t},$ and hence is a dense open subset of $\operatorname{Def}(X)$ by [Mat17, Theorem 1.2]. We will prove Theorem 4:
Theorem 2.1.
Let $X$ be a hyperkähler manifold and $L$ an isotropic nef line bundle on $X$. The locus $\operatorname{Def}(X,L)\setminus\operatorname{Def}(X,L)_{\rm nef}$ is a countable union of hypersurfaces in $\operatorname{Def}(X,L).$ Moreover, if in addition $X$ is a polarised hyperkähler manifold with ample line bundle $H$, then $\operatorname{Def}(X,L,H)_{\rm nef}$ is open and dense in $\operatorname{Def}(X,L,H)$.
This provides more evidence for the validity of the SYZ Conjectures. Here,
$\operatorname{Def}(X,L,H)_{\rm nef}\coloneqq\{t\in\operatorname{Def}(X,L,H)\mid{\mathcal{L}}_{t}\text{ is nef on }X_{t}\}.$
2.1. Wall Divisors and the Kähler cones
Denote by $\mathcal{C}_{X}\subset H^{1,1}(X)\cap H^{2}(X,{\mathbb{R}})$ the closed positive cone of $X$, i.e., the connected component of
$\{x\in H^{1,1}(X)\cap H^{2}(X,{\mathbb{R}})\mid q_{X}(x)>0\}$
that contains a Kähler class. Recall that the Kähler cone $\mathcal{K}_{X}\subset\mathcal{C}_{X}$ is the open convex cone of all Kähler classes on $X$. The birational Kähler cone $\mathcal{B}\mathcal{K}_{X}\subset\mathcal{C}_{X}$ is the union of $f^{*}\mathcal{K}_{X^{\prime}}$, where $f:X\dashrightarrow X^{\prime}$ is a birational map with $X^{\prime}$ a hyperkähler manifold. An isotropic primitive line bundle $L\in H^{1,1}(X)$ is nef if and only if $L\in\overline{{\mathcal{K}}_{X}}$, and hence we have:
$\operatorname{Def}(X,L)_{\rm nef}\coloneqq\{t\in\operatorname{Def}(X,L)\mid{\mathcal{L}}_{t}\in\overline{\mathcal{K}_{X_{t}}}\}.$
Definition 2.2.
Let $D\in\operatorname{Pic}(X)$ be a primitive divisor, not necessarily effective. Then $D$ is a wall divisor if $q_{X}(D)<0,$ and $\gamma(D^{\perp})\cap\mathcal{B}\mathcal{K}_{X}=\varnothing$ for all $\gamma\in\operatorname{Mon}^{2}(X)_{\rm Hdg}.$
Here, $\operatorname{Mon}^{2}(X)_{\rm Hdg}$ is the subgroup of monodromy operators that preserve the Hodge structure on $H^{2}(X,\mathbb{C})$, and $D^{\perp}=\{\alpha\in\overline{\mathcal{C}_{X}}\mid q_{X}(\alpha,D)=0\}.$ Thus a line bundle $L$ is nef on $X$ provided $q_{X}(L,D)\geq 0$ for all wall divisors $D\in\operatorname{Pic}(X).$
Note that by [AV17, Cor 1.6 and discussion after], for a fixed deformation type, the Beauville-Bogomolov square of a wall divisor $D$ is bounded below, i.e., there exists $N\in\mathbb{N}$ such that $-N\leq q_{Y}(D)<0$. We thus define:
$\mathcal{W}:=\{v\in\Lambda\mid-N\leq v^{2}<0,v\text{ primitive }\}.$
For any $t\in\operatorname{Def}(X)$ and any wall divisor $D_{t}\in H^{1,1}(X_{t}),$ we have $\varphi(D_{t})\in\mathcal{W}.$
The notion of wall divisors is equivalent to that of MBM classes [AV20]. Indeed, via the BBF form $q_{X}$, we can embed $H_{2}(X,{\mathbb{Z}})\subset H^{2}(X,{\mathbb{Q}})$ as an overlattice of $H^{2}(X,{\mathbb{Z}})\cong\Lambda$. We can then view classes of curves in $X$ as rational $(1,1)$-classes. Note that if $D\in\operatorname{Pic}(X)$ is a wall divisor which is positive on a Kähler class, and satisfies $D^{\perp}\cap\overline{\mathcal{K}_{X}}\neq\varnothing$, then $D$ has a rational multiple represented by an extremal curve class - an MBM class [Mon15, Prop 1.5].
2.2. Proof of Theorem 4
The proof follows from the two propositions below.
Proposition 2.3.
In the notation of Theorem 4, the locus $\operatorname{Def}(X,L)\setminus\operatorname{Def}(X,L)_{\operatorname{nef}}$ is a countable union of hypersurfaces in $\operatorname{Def}(X,L).$
Proof.
Let $t\in\operatorname{Def}(X,L)\setminus\operatorname{Def}(X,L)_{\operatorname{nef}}$. Since $L_{t}\coloneqq\mathcal{L}|_{X_{t}}$ is not nef on $X_{t}$, there exists a wall divisor $D_{t}\in\operatorname{Pic}(X_{t})$ such that $D_{t}^{\perp}\cap\overline{\mathcal{K}_{X_{t}}}\neq\varnothing$, i.e., the wall $D_{t}^{\perp}$ is a wall of the Kähler cone, and $q_{X_{t}}(D_{t},L_{t})<0$. Then the class $D_{t}\in H_{2}(X,{\mathbb{Z}})$ generates an extremal ray of the Mori cone and some positive rational multiple $kD_{t}$ is represented by an effective rational curve $C\subset X_{t}.$
Let $v:=\varphi_{t}(D_{t})\in{\mathcal{W}}$. Recall, in the notation of §1, that
$v^{\perp}(\operatorname{Def}(X,L)):=\{s\in\operatorname{Def}(X,L)\mid\varphi_{s}^{-1}(v)\in H^{1,1}(X_{s})\cap H^{2}(X,{\mathbb{Z}})\}.$
Consider $t^{\prime}\in v^{\perp}(\operatorname{Def}(X,L))$. Since the rational curve $C$ deforms exactly to the locus where $D_{t}$ remains of type $(1,1)$, it follows that there is a curve $C_{t^{\prime}}$ (possibly non integral) on $X_{t^{\prime}}$ such that
$L_{t^{\prime}}\cdot C_{t^{\prime}}=kq_{X}(L_{t},D_{t})<0$
(see [AV20, Discussion after Defn 2.10]). It follows that $L_{t^{\prime}}$ is not nef on $X_{t^{\prime}}$, and $v^{\perp}(\operatorname{Def}(X,L))\cap\operatorname{Def}(L,X)_{\operatorname{nef}}=\varnothing.$ Thus
$\operatorname{Def}(X,L)\setminus\operatorname{Def}(X,L)_{\operatorname{nef}}=\bigcup\limits_{\begin{subarray}{c}(v\cdot\varphi(L))<0\\ v\in\mathcal{W}\end{subarray}}v^{\perp}(\operatorname{Def}(X,L)),$
a union of hypersurfaces. Since ${\mathcal{W}}$ is countable, the union is also countable. ∎
Proposition 2.4.
Let $X$ be a polarised hyperkähler manifold with ample bundle $H$. Then $\operatorname{Def}(X,L,H)_{\operatorname{nef}}$ is open and dense in $\operatorname{Def}(X,L,H).$
Proof.
By Proposition 2.3 we see that
$\operatorname{Def}(X,L,H)\setminus\operatorname{Def}(X,L,H)_{\operatorname{nef}}=\bigcup\limits_{\begin{subarray}{c}(v\cdot\varphi(L))<0\\ v\in\mathcal{W}\end{subarray}}v^{\perp}(\operatorname{Def}(X,L,H)).$
We will show that, for any $t\in\operatorname{Def}(X,L,H)_{\operatorname{nef}}$, there is an open subset $U$ containing $t$ such that $U\subset\operatorname{Def}(X,L,H)_{\operatorname{nef}}$.
Put $h:=\varphi_{t}(H)$, and denote by $\Omega_{h}\coloneqq\Omega\cap\mathbb{P}(h^{\perp}\otimes\mathbb{C})$, the associated period domain. We denote by $\operatorname{Mon}(\Lambda,h)\coloneqq\{g\in\operatorname{Mon}(X_{t})\mid g(h)=h\}$. Since the lattice $h^{\perp}\subset\Lambda$ has signature $(2,*)$, by [Huy16, Remark 6.1.10] the group $\operatorname{Mon}(\Lambda,h)$ acts properly discontinuously on $\Omega_{h}$ and hence for any $v\in h^{\perp}$, the union $\bigcup_{g\in\operatorname{Mon}(\Lambda,h)}g(v)^{\perp}$ is closed. By [AV20, Theorem 3.7], the monodromy group $\operatorname{Mon}(\Lambda,h)$ acts on the set of wall divisors ${\mathcal{W}}$ with finitely many orbits. Therefore the set
$\bigcup\limits_{\begin{subarray}{c}g\in\operatorname{Mon}(\Lambda,h)\\ v^{2}<0\\ q(v,\varphi_{t}(L_{t}))<0\end{subarray}}g(v)^{\perp}$
is closed in $\Omega_{h}$. The complement $U$ of this set in $\operatorname{Def}(X,L,H)$ is an open set containing $t$. Further, $U$ is contained in $\operatorname{Def}(X,L,H)_{\operatorname{nef}}$ since, by the universality of the Kuranishi space, we have $\operatorname{Def}(X_{t},L_{t},H_{t})=\operatorname{Def}(X,L,H)$ for $t$ sufficienly close to 00.
The density follows immediately since the complement of $\operatorname{Def}(X,L,H)_{\operatorname{nef}}$ is a countable union of complex analytic codimension one subsets. ∎
As a consequence of the theorem above, the equivalent criterion for such openness due to Moriwaki [Mor92, Theorem 1] implies the following:
Corollary 2.5.
Let ${\mathcal{X}}\to S$ be a non-isotrivial family of marked hyperkähler manifolds with a fixed polarization $H$ over a smooth base scheme $S$. Suppose that $\mathcal{L}$ is a line bundle on ${\mathcal{X}}$ such that under the marking we have $\varphi_{s}(\mathcal{L}_{s})=\alpha$ for a fixed $\alpha\in\Lambda$. Define
$\sigma_{s}\coloneqq\min\{q_{X_{s}}(H,D)\mid D\text{ is a wall divisor, }D^{\perp}\cap\overline{\mathcal{K}}_{X_{s}}\neq\varnothing\text{ and }q_{X_{s}}(D,\alpha)<0\}.$
Then there exists an integer $N$, such that for all $s\in S$, $\sigma_{s}<N$.
Proof.
Since, on hyperkähler manifolds, we can detect nefness by intersecting with wall divisors (with respect to $q_{X}$), we are able to mimic Moriwaki’s argument to obtain the result. Indeed, let $X={\mathcal{X}}_{0}$, then by [Bea83, Théorèm 5, Corollaire 1] there is a universal family over $\operatorname{Def}(X,H,L)$ and there is a map $\varphi\colon S\to\operatorname{Def}(X,H,L)$ such that ${\mathcal{X}}$ is the pullback of the universal family over $\operatorname{Def}(X,H,L)$. By Theorem 4 above, there is an open subset $U\subset S$ such that for all $s\in U$, $\mathcal{L}_{s}$ is nef on $X_{s}$. In particular, for all $s\in U$, $\sigma_{s}=0$. Let $S^{\prime}\subset S\setminus U$ be an irreducible component. We can conclude from [Mor92, Lemma 2] that for all $s\in S^{\prime}$ we have $\sigma_{s}<\sigma_{\eta_{S^{\prime}}}$ where $\eta_{S^{\prime}}$ is the generic point of $S^{\prime}$. This concludes the result.
Acknowledgments
We are grateful to BIRS in Banff, and to the organizers of the workshop “Women in Mathematical Physics II” in August 2023, where this project started. We would like to thank Ignacio Barros, Emma Brakkee, Daniel Huybrechts, and Claire Voisin for interesting conversations and comments on the results. The third named author is partially supported by a grant from the Simons Foundation/SFARI (522730, LK).